OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (-2, 2) and is parallel to y = 4x + 7.

hero (hero):

-1/4(x+2) = (y-2) x/4 - 1/2 = y - 2 x/4 -1/2 + 2 -x/4 +3/2 = y